Isopropyl alcohol, CH3CHOHCH3, is used in rubbing alcohol mixtures. Alcohol on the skin cools by evaporation. How much heat is absorbed by the alcohol if 4.39 g evaporates? The heat of vaporization of isopropyl alcohol is 42.1 kJ/mol
Answer #1

Answer

moles of isopropyl alcohol = mass/molar mass = 4.39/60.1=0.073 moles

Heat = moles*heat os vaporization of isopropyl alcohol

Heat = 0.073*42.1= 3.07 kJ